OpenStudy (anonymous):

What is the y-intercept of the graph of 4x+2y=12?

OpenStudy (magbak):

@c0decraker

OpenStudy (magbak):

@codecracker

OpenStudy (anonymous):

put this in the form of y=mx + b, and the "b" will be the y intercept value: 4x+2y=12 subtract 4x from both sides: 2y=-4x+12 Divide both sides by 2: y=-2x+6 the y intercept occurs at (0,6) @Jonathan159